# A Primal-Dual-Assisted Penalty Approach to Bilevel Optimization with Coupled Constraints

### v1

@@ F.3 Numerical results for the 9-node network
Figure 10 is the counterpart of Figure 8 for the 9-node scenario, showing the behavior of our BLOCC algorithm for $\eta=1.6e-4$ and $\gamma\in\{2,3,4\}$. Each simulation is repeated 10 times (using 10 different random initializations of the upper-level variables), and both the mean and the standard deviation values are shown. Since the number of variables and constraints is almost one order of magnitude larger, the algorithm requires more time to converge. However, convergence takes place in a reasonable amount of time (20-40 times longer than in the previous 3-node test case). Regarding the optimal value, we observe that: i) the sensitivity of the (steady-state) optimal value with respect to $\gamma$ is not too large; ii) the solutions based on $y_{F,T}^{T_{F}}$ yield better upper-level values than those based on $y_{g,T}^{T_{g}}$; and iii) the gap between $f(x_{T},y_{g,T}^{T_{g}})$ and $f(x_{T},y_{F,T}^{T_{F}})$ decreases as $\gamma$ increases. Observation ii) is due to the fact that $y_{F,T}^{T_{F}}$ is not feasible (meaning that it violates the optimality of the lower level); hence, it is able to achieve a better upper-level objective. In addition, the behavior observed in iii) is consistent with the discussion in section 2.2, with the value of $\gamma$ having an impact on the suboptimality of $y_{F,t}^{T_{F}}$ at the lower level. Specifically, higher values of $\gamma$ push $y_{F,t}^{T_{F}}$ closer to $y_{g,t}^{T_{g}}$ and, hence, decrease the lower-level optimality gap. Finally, we must note that for the solid lines (associated with $y_{g,t}^{T_{g}}$), it holds that $f(x_{t},y_{g,t}^{T_{g}})=f(x_{t},y^{*}(x_{t}))$. This implies that if we need a solution that is feasible at the lower level, then better objective values are associated with higher values of $\gamma$.
@@ 2. A2)
Following the narrative in Sections F.2 and F.3, Figure 10 presents the evolution of the upper-level objective function with time for three values of the parameter $\gamma\in\{2,3,4\}$. The stepsize value has been set to $\eta=1.6\times 10^{-4}$, and two different initializations have been considered. Regarding the behavior of the algorithm with respect to the value of $\gamma$ and the particular output chosen ($y_{T,g}^{T_{g}}$ vs. $y_{T,F}^{T_{F}}$), the findings are very similar to those in Figure 10. Namely, smaller gaps are found for larger values of $\gamma$, and if feasibility at the lower level (i.e., consistency with the user preference level) must be preserved, better objective values are achieved for larger values of $\gamma$.

### v2

@@ F.3 Numerical results for the 9-node network
Figure 10 is the counterpart of Figure 8 for the 9-node scenario, showing the behavior of our BLOCC algorithm for $\eta=1.6\times 10^{-4}$ and $\gamma\in\{2,3,4\}$. Each simulation is repeated 10 times (using 10 different random initializations of the upper-level variables), and both the mean and the standard deviation values are shown. Since the number of variables and constraints is almost one order of magnitude larger, the algorithm requires more time to converge. However, convergence takes place in a reasonable amount of time (20-40 times longer than in the previous 3-node test case). Regarding the optimal value, we observe that: i) the sensitivity of the (steady-state) optimal value with respect to $\gamma$ is not too large; ii) the solutions based on $y_{F,T}^{T_{F}}$ yield better upper-level values than those based on $y_{g,T}^{T_{g}}$; and iii) the gap between $f(x_{T},y_{g,T}^{T_{g}})$ and $f(x_{T},y_{F,T}^{T_{F}})$ decreases as $\gamma$ increases. Observation ii) is due to the fact that $y_{F,T}^{T_{F}}$ is not feasible (meaning that it violates the optimality of the lower-level); hence, it is able to achieve a better upper-level objective. In addition, the behavior observed in iii) is consistent with the discussion in Section 2.2, with the value of $\gamma$ having an impact on the suboptimality of $y_{F,t}^{T_{F}}$ at the lower-level. Specifically, higher values of $\gamma$ push $y_{F,t}^{T_{F}}$ closer to $y_{g,t}^{T_{g}}$ and, hence, decrease the lower-level optimality gap. Finally, we must note that for the solid lines (associated with $y_{g,t}^{T_{g}}$), it holds that $f(x_{t},y_{g,t}^{T_{g}})=f(x_{t},y^{*}(x_{t}))$. This implies that if we need a solution that is feasible at the lower-level, then better objective values are associated with higher values of $\gamma$.
@@ 2. A2)
Following the narrative in Sections F.2 and F.3, Figure 10 presents the evolution of the upper-level objective function with time for three values of the parameter $\gamma\in\{2,3,4\}$. The stepsize value has been set to $\eta=1.6\times 10^{-4}$, and two different initializations have been considered. Regarding the behavior of the algorithm with respect to the value of $\gamma$ and the particular output chosen ($y_{T,g}^{T_{g}}$ vs. $y_{T,F}^{T_{F}}$), the findings are very similar to those in Figure 10. Namely, smaller gaps are found for larger values of $\gamma$, and if feasibility at the lower-level (i.e., consistency with the user preference level) must be preserved, better objective values are achieved for larger values of $\gamma$.
@@
Appendix G Sensitivity Analysis
Regarding the selection and impact of hyper-parameters on the performance of BLOCC, we conducted an ablation study on various values of the two critical parameters, $\gamma$ and $\eta$, and measured their effects on the optimal value and computational time. In this section, we present the sensitivity analysis on the toy example that we introduced in Section 4.1 and the 3-node network example for the transportation network planning problem in Section 4.3.
G.1 Sensitivity analysis for the toy example
We present in Table 5 the results of lower-level optimality $\|y_{g}^{*}(x_{T})-y_{F,T}\|$ using different $\gamma$ and $\eta$ to conduct BLOCC (Algorithm 1) on the toy example in Section 4.1.
$\|y_{g}^{*}(x_{T})-y_{F,T}\|$
$\eta$ $\gamma=0.001$ $\gamma=0.01$ $\gamma=0.1$ $\gamma=1.0$
0.001 $0.028\pm 0.042$ $0.035\pm 0.076$ $0.027\pm 0.064$ $000.000\pm 0$ 0.000
($3.314\pm 0.042$) ($3.186\pm 0.076$) ($2.049\pm 0.064$) ($1.967\pm 0.000$)
0.01 $0.020\pm 0.031$ $0.020\pm 0.041$ $0.009\pm 0.019$ $000.000\pm 0$ 0.000
($2.242\pm 0.031$) ($1.575\pm 0.041$) ($1.118\pm 0.019$) ($0.585\pm 0.000$)
0.1 $0.006\pm 0.010$ $0.017\pm 0.027$ $0.011\pm 0.033$ $000.000\pm 0$ 0.000
($1.616\pm 0.010$) ($1.475\pm 0.027$) ($1.257\pm 0.033$) ($0.383\pm 0.000$)
1.0 $0.023\pm 0.019$ $0.030\pm 0.022$ $0.020\pm 0.045$ $000.000\pm 0$ 0.000
($7.118\pm 0.019$) ($4.570\pm 0.022$) ($3.477\pm 0.045$) ($1.645\pm 0.000$)
10.0 $0.034\pm 0.123$ $0.019\pm 0.014$ $0.025\pm 0.070$ $000.000\pm 0$ 0.000
($6.565\pm 0.123$) ($4.365\pm 0.014$) ($2.808\pm 0.070$) ($1.750\pm 0.000$)
Table 5: Sensitivity analysis for the hyperparameters in Section 4.1. Top line in each cell represents the optimality gap $\|y^{*}_{g}(x_{T})-y_{F,T}\|$, while the bottom line represents the time required for the algorithm to converge. Both the mean and the standard deviation are for 40 simulations.
From the table, we can draw the following empirical observations:
  • O1)
    Larger values of $\gamma$ bring the upper-level objectives closer to optimal. This is consistent with Theorem 1 which illustrated that larger $\gamma$ improves the accuracy of the lower-level optimality. Since the obtained solution $y_{F,T}$ is closer to $y_{g}^{*}(x_{T})$ for larger $\gamma$ values, the distance between the $f(x_{T},y_{F,T})$ and $f(x_{T},y^{*}_{g}(x_{T}))$ will be closer as well.
  • O2)
    For a sufficiently small fixed $\eta$, larger $\gamma$ lead to faster convergence. This implies that smaller $\eta\leq\frac{1}{l_{F,1}}$ choice due to larger $\gamma$ will not significantly dampen the convergence time. This is because a large value of $\gamma$ increases $l_{F,1}$, sharpening the function $F_{\gamma}(x)$ and its gradient will be larger for most points. Thus, the gradient update $\eta\nabla F_{\gamma}(x_{t})$ will not be very small and thus won’t make the convergence slower.
G.2 Sensitivity Analysis for the 3-node network
We present in Table 6 the results of the upper-level objective value $f(x,y)$ in network planning problem for a 3-node network, whose detailed framework is introduced in Section F.
$\-f(x,y)$
$\eta$ $\gamma=1$ $\gamma=2$ $\gamma=3$ $\gamma=4$ $\gamma=5$ $\gamma=6$ $\gamma=7$ $\gamma=8$ $\gamma=9$ $\gamma=10$
$1.6\times 10^{-4}$ $-0.9625\pm 0.76$ $-1.4274\pm 0.28$ $-1.6081\pm 0.27$ $-1.4507\pm 0.41$ $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00
$13.99\pm 6.32$ $5.43\pm 1.94$ $7.85\pm 2.92$ $13.02\pm 4.47$ $5.85\pm 1.77$ $4.51\pm 1.72$ $3.06\pm 0.97$ $2.71\pm 0.95$ $2.18\pm 0.81$ $1.77\pm 0.68$
$3.2\times 10^{-5}$ $-0.9649\pm 0.76$ $-1.4768\pm 0.22$ $-1.6081\pm 0.27$ $-1.6214\pm 0.27$ $-1.4582\pm 0.41$ $-1.3772\pm 0.44$ $-1.1278\pm 0.69$ $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00 $00.0000\pm 0$ 0.00
$51.85\pm 12.12$ $15.61\pm 6.33$ $22.56\pm 6.29$ $20.83\pm 8.59$ $22.48\pm 7.89$ $29.13\pm 10.19$ $0.56\pm 0.31$ $14.38\pm 3.63$ $7.59\pm 2.84$ $6.05\pm 2.12$
$1.6\times 10^{-5}$ $-1.0461\pm 0.66$ $-1.4744\pm 0.22$ $-1.6081\pm 0.27$ $-1.6214\pm 0.27$ $-1.4582\pm 0.41$ $-1.3772\pm 0.44$ $-1.3808\pm 0.45$ $-1.2106\pm 0.73$ $-0.6792\pm 0.59$ $00.0000\pm 0$ 0.00
$21.26\pm 12.85$ $30.82\pm 8.59$ $46.02\pm 9.89$ $35.98\pm 11.00$ $32.53\pm 9.90$ $32.37\pm 11.80$ $40.79\pm 12.98$ $40.31\pm 13.95$ $1.47\pm 0.57$ $17.27\pm 4.27$
$3.2\times 10^{-6}$ $-0.7819\pm 0.68$ $-1.4725\pm 0.22$ $-1.6081\pm 0.27$ $-1.6214\pm 0.27$ $-1.6297\pm 0.27$ $-1.6355\pm 0.27$ $-1.4671\pm 0.42$ $-1.4700\pm 0.42$ $-1.3856\pm 0.45$ $-1.3873\pm 0.45$
$888.38\pm 256.38$ $213.22\pm 33.88$ $335.63\pm 43.42$ $253.26\pm 46.92$ $210.03\pm 44.41$ $206.14\pm 48.21$ $172.50\pm 43.31$ $153.57\pm 43.90$ $138.12\pm 38.92$ $117.68\pm 34.65$
$1.6\times 10^{-6}$ $-0.9665\pm 0.63$ $-1.4723\pm 0.23$ $-1.6840\pm 0.03$ $-1.6214\pm 0.27$ $-1.6297\pm 0.27$ $-1.6355\pm 0.27$ $-1.6397\pm 0.27$ $-1.4700\pm 0.42$ $-1.4722\pm 0.42$ $-1.4741\pm 0.42$
$805.74\pm 268.20$ $377.37\pm 54.25$ $629.90\pm 65.16$ $496.31\pm 69.79$ $464.33\pm 68.94$ $418.23\pm 67.45$ $390.74\pm 65.78$ $356.06\pm 60.70$ $334.36\pm 61.03$ $305.70\pm 57.01$
Table 6: Sensitivity analysis for hyperparameters $\eta$ and $\gamma$ in the experiment of the 3-node network transportation design described in Section 4.3 and Appendix F. Top line in each cell represents the upper-level objective value $f(x_{T},y_{g,T})$, while the bottom line represents convergence time. Both the mean and the standard deviation of 10 simulations are provided.
We know from Section F that the goal is to minimize $f(x,y)$, and negative values of $f(x,y)$ are expected. Therefore, $f(x_{T},y_{T})=0$ in the table indicates a failure to converge and we can see that
  • O3)
    Smaller $\eta$ is needed to satisfy $\eta\leq\frac{1}{l_{F,1}}$ to ensure convergence if $\gamma$ is chosen larger. The algorithm tends to converge for different $\gamma$ values when $\eta$ is small enough. This is because the Lipschitz smoothness constant $l_{F,1}$ of $F_{\gamma}(x)$ increases with $\gamma$ (Lemma 3), and the BLOCC algorithm is guaranteed to converge if $\eta\leq\frac{1}{l_{F,1}}$ (Theorem 2).
In summary, the experimental results align with the theoretical analysis, demonstrating that the penalty constant $\gamma$ is robust and can be effectively set around 10, and the stepsize should be adjusted to ensure smooth and monotonic convergence.
Appendix H Analysis of the Computational Complexity
H.1 Complexity comparison
We present in the following the computational complexity of our BLOCC algorithm in comparison with LV-HBA [75] and GAM [72] as baselines.
Method Iteration Cost Computational Cost per Iteration
BLOCC Outer loop (Algorithm 1): General setting (Theorem 3): $\tilde{\mathcal{O}}(d_{c}d_{x}+d_{x}^{2}+\epsilon^{-1}(d_{c}d_{y}+d_{y}^{2}))$;
$T=\mathcal{O}(\epsilon^{-1.5})$ Special case (Theorem 4): $\tilde{\mathcal{O}}(d_{x}^{2}+d_{y}^{2}+d_{c}(d_{x}+d_{y}))$
LV-HBA $\mathcal{O}(\epsilon^{-3})$ $\mathcal{O}(d_{x}d_{y}d_{c}+(d_{x}+d_{y})^{3.5})$
GAM Asymptotic More than $\mathcal{O}(d_{y}^{3}+d_{x}^{2}+d_{c}(d_{x}+d_{y}))$
Table 7: Complexity comparison of our work with LV-HBA [75] and GAM [72].
The iteration costs are detailed in the earlier sections, and in [75] and [72]. Hence, we discuss next the per-iteration cost. In the following discussion, we assume
  • A1)
    the complexity of calculating a function is proportional to the dimension of the inputs. For example, The complexity of finding $\nabla_{x}g(x,y),\nabla_{x}f(x,y)$ are $\mathcal{O}(d_{x})$, and the one for $g^{c}(x,y)\rangle$ is $\mathcal{O}(d_{x}d_{y})$.
  • A2)
    Projection cost on $\mathcal{X}$ and $\mathcal{Y}$ is respectively no more than $\mathcal{O}(d_{x}^{2})$ and $\mathcal{O}(d_{y}^{2})$. As discussed in the introduction, $\mathcal{X}$ and $\mathcal{Y}$ are assumed to be easy-to-project domains, i.e. projection can be done using a projection matrix or a simple formula, and the projection costs are no more than $\mathcal{O}(d_{x}^{2})$ and $\mathcal{O}(d_{y}^{2})$, respectively.
In this way, BLOCC is a first-order method with gradient calculation costs of $\mathcal{O}(d_{x}d_{c})$ and $\mathcal{O}(d_{y}d_{c})$, and the projection cost of $\mathcal{O}(d_{x}^{2})$ and $\mathcal{O}(d_{y}^{2})$. We present the detailed analysis of achieving the computational cost in the table in the following Section H.2.
In the SVM model training and network planning experiments in Section 4, projections are simpler (e.g., truncation), resulting in $\mathcal{O}(d_{x})$ and $\mathcal{O}(d_{y})$ costs. In this scenario, for the general setting (Theorem 3), the complexity is $\tilde{\mathcal{O}}(\epsilon^{-1.5}d_{c}d_{x}+\epsilon^{-2.5}d_{c}d_{y})$; and for the $g^{c}$ affine in $y$ setting (Theorem 4), it is $\tilde{\mathcal{O}}(\epsilon^{-1.5}d_{c}(d_{x}+d_{y}))$. Therefore, our BLOCC is especially robust to large-scale problems.
LV-HBA [75], also a first-order method, has similar gradient calculation costs. However, its projection onto $\{\mathcal{X}\times\mathcal{Y}:g^{c}(x,y)\leq 0\}$ is expensive, with a complexity of $\mathcal{O}((d_{x}+d_{y})^{3.5})$ of using interior point method to find the projected point [39], and evaluating $g^{c}(x,y)\leq 0$ adds $\mathcal{O}(d_{x}d_{y}d_{c})$ as it requires $d_{c}$ inequality judgment on functions taking input dimension $d_{x},d_{y}$.
GAM [72] lacks an explicit algorithm for lower-level optimality and Lagrange multipliers. Even if we omit this, calculating $\nabla_{yy}^{2}g$ and its inverse incurs a cost of $\mathcal{O}(d_{y}^{3})$. Additionally, it has cost $\mathcal{O}(d_{x}^{2}+d_{c}(d_{x}+d_{y}))$ for projection onto $\mathcal{X}$ and calculating gradients.
We can see that BLOCC stands out with the lowest computational cost in both iterational and overall complexity thanks to its first-order and joint-projection-free nature. Consequently, BLOCC is particularly well-suited for large-scale applications with high-dimensional parameters.
H.2 Complexity analysis of BLOCC
At each iteration $t$, our proposed BLOCC algorithm in Algorithm 1 involves:
  • Step 1:
    Solve two max-min problems.
  • Step 2:
    Calculate $g_{F,t}$ in (13) and update $x_{t+1}=\operatorname{Proj}_{\mathcal{X}}\big{(}x_{t}-\eta g_{F,t}\big{)}$.
where Step 1 can be achieved by our proposed max-min solver in Algorithm 2. In the following, we provide analysis for using the accelerated version of the Algorithm 2 in the general case (Theorem 3) and the single-loop version for the special case (Theorem 4) as discussed in Section 3.3.
In the general case, we use the accelerated version of Algorithm 2 to solve both (7) and (5). For solving (7), at each inner loop iteration $t$, line 3 of the algorithm is of computational cost $\mathcal{O}(d_{c})$. The update of $y$ in line 4-6 involves calculating $\nabla_{y}g(x,y)$ with a cost of $\mathcal{O}(d_{y})$, finding $\langle\mu,\nabla_{y}g^{c}(x,y)\rangle$ for fixed $x$ of $\mathcal{O}(d_{y}d_{c})$ following assumption A1), and the projection $\operatorname{Proj}_{\mathcal{Y}}$ of complexity $\mathcal{O}(d_{y}^{2})$ according to A2). Moreover, $y$ converges linearly as $L(\mu,y)$ is strongly convex in $y$ (Lemma 7), this inner update for $y$ gives an iteration complexity of $\mathcal{O}(\ln(\epsilon^{-1}))$. Therefore, the cost of the update for $y$ totals up to $\mathcal{O}(\ln(\epsilon^{-1})(d_{y}d_{c}+d_{y}^{2}))$. The update of $\mu$ in line 7 involves $\mathcal{O}(d_{y}d_{c})$ for calculating $\nabla_{\mu}L(\mu_{t+1/2},y_{t+1})=g^{c}(x,y_{t+1})$ with $x$ being fixed, and $\mathcal{O}(d_{c})$ for projection onto $\mathbb{R}_{+}^{d_{c}}$.
Moreover, to achieve target accuracy $\epsilon$ on the metric $\frac{1}{T}\sum_{t=0}^{T}\|G_{\eta}(x_{t})\|^{2}\leq\epsilon$, with $\gamma=\mathcal{O}(\epsilon^{-1})$ using the BLOCC algorithm, the iteration complexity of the max-min solver is $\mathcal{O}(\epsilon^{-1})$ according to Theorem 3. Therefore, the complexity for solving (7) in the general case using the accelerated version of Algorithm 2 (Theorem 3) is
$\mathcal{O}(\epsilon^{-1}(\ln(\epsilon^{-1})(d_{y}d_{c}+d_{y}^{2})+d_{c}d_{y}+d_{c}))=\tilde{\mathcal{O}}(\epsilon^{-1}(d_{y}d_{c}+d_{y}^{2})).$ (67)
Solving (5) is of the same order as the only difference is in calculating $\nabla_{y}f(x,y)$. In this way, we can conclude that Step 1 is at a cost of $\tilde{\mathcal{O}}(\epsilon^{-1}(d_{y}d_{c}+d_{y}^{2}))$ in the general case.
In the special case (Theorem 4), the non-accelerated (single-loop) version of Algorithm 2 involves $y_{t+1}=\operatorname{Proj}_{\mathcal{Y}}(y_{t}-\eta_{1}\nabla_{y}L(\mu_{t},y_{t}))$ with complexity $\mathcal{O}(d_{y}^{2}+d_{y}d_{c})$, and $\mu_{t+1}=\operatorname{Proj}_{\mathbb{R}_{+}^{d_{c}}}(\mu_{t}+\eta_{2}\nabla_{\mu}L(\mu_{t},y_{t+1}))$ with complexity $\mathcal{O}(d_{c}+d_{y}d_{c})$ following a similar analysis as in the general case.
Moreover, the algorithm converges linearly in the special case of $g^{c}$ being affine in $y$ and the computational complexity is $\mathcal{O}(\ln(\epsilon^{-1}))$, according to Theorem 4. Therefore, the complexity for the max-min Step 1 in the special case is
$\mathcal{O}(\ln(\epsilon^{-1})(d_{y}d_{c}+d_{y}^{2}+d_{c}))=\tilde{\mathcal{O}}(d_{y}d_{c}+d_{y}^{2}+d_{c}).$ (68)
Step 2 involves calculating $\nabla_{x}g(x,y)$ with with a fixed $y$, and $\langle\mu,\nabla_{x}g^{c}(x,y)\rangle$ for fixed $\mu,y$, which are of complexity $\mathcal{O}(d_{x})$ and $\mathcal{O}(d_{x}d_{c})$ respectively, and conducting a projection on $\mathcal{X}$ of cost $\mathcal{O}(d_{x}^{2})$ according to assumption A2).
In this way, the computational complexity of Step 2 is
$\mathcal{O}(d_{x}d_{c}+d_{x}^{2}).$ (69)
We can therefore conclude the complexity of BLOCC in Table 7.